# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v1

@@ Abstract
Computing the polar decomposition and the related matrix sign function, has been a well-studied problem in numerical analysis for decades. More recently, it has emerged as an important subroutine in deep learning, particularly within the Muon optimization framework. However, the requirements in this setting differ significantly from those of traditional numerical analysis. In deep learning, methods must be highly efficient and GPU-compatible, but high accuracy is often unnecessary. As a result, classical algorithms like Newton–Schulz (which suffers from slow initial convergence) and methods based on rational functions (which rely on QR decompositions or matrix inverses) are poorly suited to this context. In this work, we introduce Polar Express, a GPU-friendly algorithm for computing the polar decomposition. Like classical polynomial methods such as Newton–Schulz, our approach uses only matrix-matrix multiplications, making it GPU-compatible. Motivated by earlier work of Chen & Chow and Nakatsukasa & Freund, Polar Express adapts the polynomial update rule at each iteration by solving a minimax optimization problem, and we prove that it enjoys a strong worst-case optimality guarantee. This property ensures both rapid early convergence and fast asymptotic convergence. We also address finite-precision issues, making it stable in bfloat16 in practice. We apply Polar Express within the Muon optimization framework and show consistent improvements in validation loss on large-scale models such as GPT-2, outperforming recent alternatives across a range of learning rates.
@@ 1   Introduction
Advanced linear algebra is making its way into deep learning. In particular, efficient algorithms for computing matrix functions have found applications in training neural networks: approximations to the matrix-inverse are used in the full Adagrad method [14], the matrix square-root and quarter-root appear as subroutines in the Shampoo optimizer [18, 42], and recently, the matrix sign function has found applications within the Muon-optimizer [6, 5, 23].
While the problem of computing these matrix functions has been studied by numerical analysts for decades, applications in deep learning come with different requirements than those in computational science. For example, it is critical to develop methods that take advantage of GPU-friendly operations like matrix-matrix products and avoid less parallel operations. Moreover, for large models, limiting memory overhead is critical. On the other hand, accuracy is typically less critical in machine learning — the long-standing gold standard of $10^{-16}$ double precision is overkill in deep learning.
With these new considerations in mind, there is an opportunity for research on tailoring matrix function algorithms specifically for deep learning [2]. In this paper, we take a step in this direction, engineering a state-of-the-art, GPU-friendly algorithm for computing the matrix sign function, or more generally, for computing the polar decompostion of a rectangular gradient matrix for updating neural network weights. We apply our new Polar Express method to computing the descent direction in the recently popular Muon method. In Figure 1 we show how using our Polar Express method, when coupled with the Muon method, consistently achieves a lower validation loss of GPT-2 model for all learning rates, as compared to recent alternative methods [10, 43, 23].
@@ 1.1   The Muon Method
Muon has recently gained popularity for training large language models, often outperforming state-of-the-art adaptive gradient methods like Adam and AdamW [27, 32]. Muon has been used to set records for the NanoGPT speedrun [23], and to expand the Pareto frontier for efficient training of large language models [31, 41].
Figure 1: Training a GPT-2 (124M) model on 1 Billion tokens of Fineweb data set [3]. The Legend muon-<name> refers to using muon with the <name> method for computing the polar. Left: The final Validation loss vs the learning rate and Right: Validation vs iterations for each method with the best resulting learning rate. The final best validation loss for each method, in reverse order was adamw: $4.197$, muon-Jordan: $3.639$, muon-You: $3.629$ and muon-PolarExp: $3.588$.
The Muon update rule [6] is defined as follows. Let $\lambda,\beta>0$ be the learning rate and momentum coefficient hyperparameters. (By default, $\beta=0.9$.) Let $\bm{W}_{t}\in\mathbb{R}^{m\times n}$ be the weight matrix of a given neural network layer at iteration $t$, and let $\bm{G}_{t}\in\mathbb{R}^{m\times n}$ be its (stochastic) gradient. Let $\bm{M}_{t}\in\mathbb{R}^{m\times n}$ be the momentum estimate of the gradient, where $\bm{M}_{0}=\bm{0}$. The Muon update is given by
@@
While standard stochastic gradient descent (SGD) with momentum updates the weight matrix by taking a step in the direction $-\bm{M}_{t}$, the Muon method steps in the direction $-\operatorname*{polar}(\bm{M}_{t})$, where $\operatorname*{polar}(\bm{M})$ denotes the closest semi-orthogonal matrix to $\bm{M}$ [20, Chapter 8]. Concretely, if $\bm{M}=\bm{U}\bm{\Sigma}\bm{V}^{\mathsf{T}}$ is the rank-reduced singular value decomposition (SVD) of $\bm{M}$, then
@@
The matrix $\operatorname*{polar}(\bm{M})$ can be seen as a generalization of the matrix sign function to rectangular matrices [4]. Indeed, when $\bm{M}$ is square symmetric with eigendecomposition $\bm{M}=\bm{V}\bm{\Lambda}\bm{V}^{\mathsf{T}}$, $\operatorname*{polar}(\bm{M})$ exactly coincides with the matrix sign function $\operatorname*{sign}(\bm{M})=\bm{V}\operatorname*{sign}(\bm{\Lambda})\bm{V}^{\mathsf{T}}$ [20, Chapter 10]. Equivalently, $\operatorname*{polar}(\bm{M})$ is the left orthogonal factor of the polar decomposition of $\bm{M}$ [20, Chapter 8]. The motivation for Muon is that the direction $-\operatorname*{polar}(\bm{M})$ corresponds to the steepest-descent step with respect to the spectral norm of the weight matrix $\bm{W}_{t}$ (instead of the Frobenius norm, as in standard SGD). We refer the reader to [23] or [6] for more information. In this paper, we take the Muon update rule for granted and study the problem of computing $\operatorname*{polar}(\bm{M})$ efficiently.
@@ 1.2   Computing the Polar Factor
Although $\operatorname*{polar}(\bm{M})$ can be computed directly via an SVD in $O(\min(mn^{2},nm^{2}))$ time, doing so is prohibitively expensive in deep learning applications, especially as standard SVD algorithms fail to take full advantage of the parallelism available on GPUs. There has been significant work on highly-parallel methods for the SVD, but the most common approaches compute the matrix-sign function or polar decomposition as a subroutine [34, 36]. Numerical analysts have spent decades developing iterative methods for computing $\operatorname*{polar}(\bm{M})$. This rich line of work includes Newton–Schulz [20, Chapter 8], Padé iteration [26, 19], the Newton and scaled Newton iterations [20, Chapter 8], the QWHD iteration [33, 36], and Zolo-pd [34]. Unfortunately, as outlined in Section 2, most of these methods are based on rational approximations to the function $\operatorname*{sign}(x)$ and require computing matrix inverses or QR decompositions. Such methods are ill-suited to GPU acceleration and deep learning applications. In contrast, the older Newton-Schulz method is based on polynomial approximation of $\operatorname*{sign}(x)$ and uses only matrix-matrix products. Thus, Muon initially used Newton-Schulz [5]. Indeed, Muon stands for “MomentUm Orthogonalized by Newton-Schulz” [23].
The Newton-Schulz method constructs a sequence of approximations $\bm{X}_{t}\approx\operatorname*{polar}(\bm{M})$ as follows:
@@
At each iteration, this rule effectively applies the cubic polynomial $p(x)=\frac{3}{2}x-\frac{1}{2}x^{3}$ to the singular values of $\bm{X}_{t}$. It is well-known that the scalar fixed point iteration $x_{t+1}=p(x_{t})$ converges to $\operatorname*{sign}(x_{0})$ as $t\to\infty$, provided $|x_{0}|\leq 1$. As a result, the matrix iteration satisfies $\lim\limits_{t\to\infty}\bm{X}_{t}=\bm{U}\bm{V}^{\top}=\operatorname*{polar}(\bm{X}_{0})$. Higher-degree versions of Newton-Schulz follow the same principle. For example, the degree 5 polynomial $p(x)=(15x-10x^{3}+3x^{5})/8$ can be used to accelerate the convergence. While the Newton-Schulz iteration converges super-exponentially when $\bm{X}_{t}$ is sufficiently close to $\operatorname*{polar}(\bm{M})$, it suffers from slow initial convergence; when $\bm{X}_{0}$ is far from $\operatorname*{polar}(\bm{M})$, the approximation improves slowly over the first few iterations.
In Muon, high accuracy approximations to $\operatorname*{polar}(\bm{M})$ are usually not necessary. The primary goal is instead to compute a coarse approximation in as few iterations as possible. To accelerate convergence in the low-accuracy regime, Jordan recently proposed a fixed-point iteration based on the polynomial $p(x)=3.4445x-4.7750x^{3}+2.0315x^{5}$ [23], which was found using a heuristic numerical search. Unlike Newton-Schulz, the scheme that Jordan proposed does not converge to $\operatorname*{polar}(\bm{M})$. Instead, it plateaus at an error of $\approx 0.3$. However, it reaches this level of accuracy rapidly. As a result, when the number of iterations is smaller than $10$, Jordan’s method outperforms the Newton-Schulz iteration. Building on this idea, You [10] proposed a method that applies six different polynomial updates in succession. This method can achieve somewhat better accuracy than Jordan’s but still fails to converge.
@@ 1.3   Contributions
We present Polar Express, an iterative method for approximating $\operatorname*{polar}(\bm{M})$. Our method dynamically adapts the polynomial update rule at each iteration, prioritizing rapid progress in the initial stage and high accuracy in the later stage. Polar Express constructs polynomials $p_{1},\ldots,p_{T}$ so that the resulting composition is the optimal approximation to the sign function with respect to the supremum ($L^{\infty}$) norm (Theorem 4.1). By iteratively applying these polynomials to $\bm{M}$, Polar Express inherits a worst-case optimality guarantee for approximating $\operatorname*{polar}(\bm{M})$ at every iteration. The method converges to $\operatorname*{polar}(\bm{M})$ super-exponentially (Theorem 4.3), and it quickly reaches a reasonable approximation in just five or ten iterations. This early-stage acceleration is especially valuable in deep learning applications, where runtime efficiency takes precedence over high accuracy. In contrast, classical methods like Newton-Schulz suffer from a slow initial convergence, while recent heuristic proposals [23, 10] fail to converge. Our method is efficient to run on GPUs, using only a few matrix-matrix products per iteration.
We give an explicit instantiation of Polar Express in Section 4.4 with minor modifications to make it compatible with half-precision arithmetic. This formula can be used as a drop-in replacement for previous methods. In numerical experiments, our method outperforms previous methods on synthetic matrices and gradient matrices from a GPT-2 transformer (Fig 4). We demonstrate the effectiveness of Polar Express within the Muon algorithm (Figure 1), showing that it consistently improves the training of GPT-2 language models on 1 billion tokens of the Fineweb dataset [3].
@@
For a matrix $\bm{M}$ with rank reduced SVD given by $\bm{M}=\bm{U}\bm{\Sigma}\bm{V}^{\mathsf{T}}$ and positive singular values $\sigma_{1}\geq\cdots\geq\sigma_{\operatorname*{rank}(\bm{M})}>0$, and a polynomial $p\in\mathbb{P}_{d}^{\operatorname*{odd}}$, we denote by $p(\bm{M})=\bm{U}p(\bm{\Sigma})\bm{V}^{\mathsf{T}}$, where $p(\bm{\Sigma})$ is the diagonal matrix with diagonal entries $p(\sigma_{i})$ for $i=1,\ldots,\operatorname*{rank}(\bm{M})$.
@@ 2   Related Work
The earliest methods in the literature are polynomial iterations like (2). Several nearly simultaneous papers introduced the family of polynomial Padé iterations, comprising Newton-Schulz and its higher-degree analogues [28, 7, 19, 30]. These higher-degree methods are also sometimes called “Newton-Schulz”; when doing so, we will specify the degree for clarity. In these methods, each iteration refines the current approximation $\bm{X}_{t}$ by applying a low-degree odd matrix polynomial, where any odd monomial $x\mapsto x^{2q+1}$ is defined for rectangular matrices by the formula $\bm{X}_{t}\mapsto\bm{X}_{t}\left(\bm{X}_{t}^{\top}\bm{X}_{t}\right)^{q}$. Our Polar Express method also takes this form, though it changes the polynomial at each iteration. The polynomials used in Padé methods are chosen to match the value and first few derivatives of $\operatorname*{sign}(x)$ at the points $x=\pm 1$. For instance, the update rule of the third method in this family is defined by $p(x)=\frac{1}{16}\left(35x-35x^{3}+21x^{5}-5x^{7}\right)$, which is the unique degree-7 polynomial satisfying $p(\pm 1)=1$ and $p^{\prime}(\pm 1)=p^{\prime\prime}(\pm 1)=p^{\prime\prime\prime}(\pm 1)=0$. These methods converge so long as all singular values of $\bm{X}_{0}$ lie in $(0,1]$, a condition guaranteed by the initialization of (2). Furthermore, the order of convergence of the degree $12q+$ 1 method is $q+1$ [7]. In particular, the Newton-Schulz method ($q=1$) converges quadratically.
In the numerical analysis literature, these polynomial methods were succeeded by rational iterations like Newton’s method [19], defined as follows [^1] We describe Newton’s method and other rational methods for square non-singular $\bm{M}$. Non-square problems can be reduced to the square case by an initial QR decomposition, but this is not an option for purely polynomial methods like ours.:
@@
Newton’s method also converges quadratically, but it uses fewer operations per iteration than Newton-Schulz. Like Newton-Schulz, it works because the rational function $r(x)=\frac{1}{2}(x+x^{-1})$ has a stable fixed point at $1$; unlike for Newton-Schulz, this point is a global attractor for the whole positive real line. At first glance, Newton’s method has nothing to do with the Padé iterations discussed above. However, after a change of variables $\bm{Y}_{t}=\bm{X}_{t}^{-1}$, it can be reinterpreted as $\bm{Y}_{t+1}=2\bm{Y}_{t}(\bm{I}+\bm{Y}_{t}^{\top}\bm{Y}_{t})^{-1}$, which is sometimes called inverse Newton. Observing that $r(x)=\frac{2x}{1+x^{2}}$ satisfies $r(\pm 1)=1$ and $r^{\prime}(\pm 1)=0$, we see that (inverse) Newton is also a Padé method, though a rational rather than polynomial one. In fact, given a odd degree $12q_{n}+$ 1 for the numerator and an even degree $2q_{d}$ for the denominator, there is a unique rational function that matches the value and first $q_{n}+q_{d}$ derivatives of $\operatorname*{sign}(x)$ at $x=\pm 1$. This directly yields a Padé method for computing $\operatorname*{polar}(\bm{M})$ whose order of convergence is $q_{n}+q_{d}+1$. For instance, $r(x)=\frac{x(3+x^{2})}{1+3x^{2}}$ is called Halley’s method, which converges cubically. When $q_{d}=0$, we recover the polynomial Padé methods.
@@
There are two main weakness of Newton’s method and the Padé iterations: slow convergence in the initial phase and the need to compute explicit inverses. To accelerate initial convergence, Higham popularized the technique of rescaling the matrix after every Newton iteration [19]. Intuitively, rescaling $\bm{X}_{t}$ so that $\sigma_{\max}=1/\sigma_{\min}$ centers the spectrum around $1$, where convergence is fastest. Since the (inverse) Newton update treats inputs in the interval $(1/z,1)$ exactly like those in in $(1,z)$, this recentering can only help. Several easily-computable choices of scaling factor exist to accomplish this approximately. Computing matrix inverses is difficult to parallelize and to implement stably in low precision arithmetic. However, a trick was developed for stably computing many rational methods without explicit inverses; QR decompositions can be used instead [33, 47]. Applying this trick to Halley’s method and combining with a special rescaling scheme yields the QR-based dynamically weighted Halley iteration (QDWH), which converges in just six iterations for any reasonably conditioned matrix [33].

### v2

@@ Abstract
Computing the polar decomposition and the related matrix sign function has been a well-studied problem in numerical analysis for decades. Recently, it has emerged as an important subroutine within the Muon algorithm for training deep neural networks. However, the requirements of this application differ sharply from classical settings: deep learning demands GPU-friendly algorithms that prioritize high throughput over high precision. We introduce Polar Express, a new method for computing the polar decomposition. Like Newton–Schulz and other classical polynomial methods, our approach uses only matrix-matrix multiplications, making it very efficient on GPUs. Inspired by earlier work of Chen & Chow and Nakatsukasa & Freund, Polar Express adapts the update rule at each iteration by solving a minimax optimization problem. We prove that this strategy minimizes error in a worst-case sense, allowing Polar Express to converge as rapidly as possible both in the early iterations and asymptotically. We also address finite-precision issues, making it practical to use in bfloat16. When integrated into the Muon training framework, our method leads to consistent improvements in validation loss when training a GPT-2 model on one billion tokens from the FineWeb dataset, outperforming recent alternatives across a range of learning rates.
@@ 1   Introduction
Advanced linear algebra is making its way into deep learning. Efficient algorithms for computing matrix functions have found exciting new applications in training neural networks. In particular, approximations to the matrix-inverse are used in the full Adagrad method [13], the matrix square-root and quarter-root appear as subroutines in the Shampoo and Soap optimizers [17, 43, 47], and most recently, the matrix sign function has become a key ingredient of the Muon optimizer [5, 4, 22].
While the problem of computing these matrix functions has been studied by numerical analysts for decades, applications in deep learning come with different requirements than those in computational science. For deep learning, it is critical to take maximum advantage of GPU-friendly operations like matrix-matrix products and to avoid less parallel operations. Moreover, memory overhead must be small to handle large models. On the other hand, high accuracy is typically less important; the gold standard of sixteen digits of accuracy is overkill in deep learning.
Given these considerations, there is a need to develop new matrix function methods that are tailor-made for deep learning applications. We take on this challenge by designing a state-of-the-art, GPU-friendly algorithm for computing the matrix sign function, or more generally, for computing the polar decomposition of a rectangular matrix. We apply our new Polar Express method (Algorithm 1) to compute the descent direction in the increasingly popular Muon optimizer. In Figure 1, we show that using Polar Express within Muon consistently results in lower validation loss across all learning rates when training a GPT-2 model, as compared to other matrix sign methods [9, 44, 22].
Algorithm 1 Python code for the Polar Express of degree = 5.
from itertools import repeat
import torch
coeffs_list = [
(8.28721201814563, -23.595886519098837, 17.300387312530933),
(4.107059111542203, -2.9478499167379106, 0.5448431082926601),
(3.9486908534822946, -2.908902115962949, 0.5518191394370137),
(3.3184196573706015, -2.488488024314874, 0.51004894012372),
(2.300652019954817, -1.6689039845747493, 0.4188073119525673),
(1.891301407787398, -1.2679958271945868, 0.37680408948524835),
(1.8750014808534479, -1.2500016453999487, 0.3750001645474248),
(1.875, -1.25, 0.375), # subsequent coeffs equal this numerically
# safety factor for numerical stability (but exclude last polynomial)
coeffs_list = [(a / 1.01, b / 1.01**3, c / 1.01**5)
for (a, b, c) in coeffs_list[:-1]] + [coeffs_list[-1]]
@torch.compile
def PolarExpress(G: torch.Tensor, steps: int) -> torch.Tensor:
assert G.ndim >= 2
X = G.bfloat16() # for speed
if G.size(-2) > G.size(-1): X = X.mT # this reduces FLOPs
X = X / (X.norm(dim=(-2, -1), keepdim=True) * 1.01)
hs = coeffs_list[:steps] + list(
repeat(coeffs_list[-1], steps - len(coeffs_list)))
for a, b, c in hs:
A = X @ X.mT
B = b * A + c * A @ A
X = a * X + B @ X # X <- aX + bX^3 + cX^5
if G.size(-2) > G.size(-1): X = X.mT
return X
Figure 1: Training a GPT-2-Large model (774M params) on 1 billion tokens from the FineWeb dataset [2]. The label muon-<name> refers to implementing Muon using <name> to compute the polar factor. Left: final validation loss across learning rates. Right: validation loss across epochs using the best learning rate. The best learning rate ($lr$) and final validation loss for each method was adamw $(lr=0.0001)$: $4.172$, muon-You $(lr=0.02)$: $3.400$, muon-Jordan $(lr=0.02)$: $3.398$ and muon-PolarExp $(lr=0.02)$: $3.340$.
@@ 1.1   The Muon Method
The Muon optimizer has recently gained popularity for training large language models, often outperforming state-of-the-art adaptive gradient methods like Adam and AdamW [26, 31]. Muon has been used to set records for the NanoGPT speedrun [22] and to expand the Pareto frontier of performance versus training FLOPs for large language models [30, 42].
The Muon update rule [5] is defined as follows. Let $\lambda,\beta>0$ be the learning rate and momentum coefficient hyperparameters. (By default, $\beta=0.9$.) Let $\bm{W}_{t}\in\mathbb{R}^{m\times n}$ be the weight matrix of a given neural network layer at iteration $t$, and let $\bm{G}_{t}\in\mathbb{R}^{m\times n}$ be its (stochastic) gradient. Let $\bm{M}_{t}\in\mathbb{R}^{m\times n}$ be the running momentum estimate of the gradient, where $\bm{M}_{0}=\bm{0}$. The Muon update is given by
@@
Whereas standard stochastic gradient descent (SGD) with momentum updates the weight matrix by taking a step in the direction $-\bm{M}_{t}$, the Muon method steps in the direction $-\operatorname*{polar}(\bm{M}_{t})$, where $\operatorname*{polar}(\bm{M})$ denotes the closest semi-orthogonal matrix to $\bm{M}$ [19, Chapter 8]. Concretely, if $\bm{M}=\bm{U}\bm{\Sigma}\bm{V}^{\mathsf{T}}$ is the singular value decomposition (SVD) of $\bm{M}$, then
@@
The matrix $\operatorname*{polar}(\bm{M})$ can be seen as a generalization of the matrix sign function to rectangular matrices [3]. Indeed, when $\bm{M}$ is square symmetric with eigendecomposition $\bm{M}=\bm{V}\bm{\Lambda}\bm{V}^{\mathsf{T}}$, $\operatorname*{polar}(\bm{M})$ exactly coincides with the matrix sign function $\operatorname*{sign}(\bm{M})=\bm{V}\operatorname*{sign}(\bm{\Lambda})\bm{V}^{\mathsf{T}}$ [19, Chapter 5]. Equivalently, $\operatorname*{polar}(\bm{M})$ is the left orthogonal factor of the polar decomposition of $\bm{M}$ [19, Chapter 8]. The motivation for Muon is that $-\operatorname*{polar}(\bm{M})$ gives the steepest-descent direction with respect to the spectral norm (instead of the Frobenius norm, as in standard SGD).
Recent work [40] shows that Muon can be viewed as a conditional gradient (Frank-Wolfe) method with a trust region defined by the spectral norm. In the same work, the authors also provide a convergence theory for the smooth and non-convex setting, as well as for the stochastic non-convex case. The analysis of Muon was further refined in [41], which proves convergence under a layerwise $(L_{0},L_{1})$-smoothness assumption, in both the stochastic non-convex and stochastic Polyak–Łojasiewicz settings. We refer the reader to [22] and [5] for further background. In this paper, we take the Muon update rule as given and focus on the problem of efficiently computing the polar decomposition $\operatorname*{polar}(\bm{M})$.
@@ 1.2   Computing the Polar Factor
Although $\operatorname*{polar}(\bm{M})$ can be computed directly via an SVD in $O(\min(mn^{2},nm^{2}))$ time, doing so is prohibitively expensive in deep learning applications, especially as standard SVD algorithms fail to take full advantage of the parallelism available on GPUs. There has been significant work on highly-parallel methods for the SVD, but the most common approaches actually require computing the matrix-sign function as a subroutine [33, 35]. Numerical analysts have spent decades developing iterative methods for computing $\operatorname*{polar}(\bm{M})$. This rich line of work includes Newton–Schulz [19, Chapter 8], Padé iteration [25, 18], the Newton and scaled Newton iterations [19, Chapter 8], the QWHD iteration [32, 35], and Zolo-pd (Zolotarev polar decomposition) [33]. Unfortunately, as discussed in Section 2, most of these methods are based on rational approximations to the function $\operatorname*{sign}(x)$ and require computing matrix inverses or QR decompositions. Such methods are ill-suited to GPU acceleration and deep learning applications. In contrast, the older Newton-Schulz method is based on polynomial approximation of $\operatorname*{sign}(x)$ and uses only matrix-matrix products. Thus, Muon initially used Newton-Schulz [4]. Indeed, Muon stands for “MomentUm Orthogonalized by Newton-Schulz” [22].
The Newton-Schulz methods.
Newton-Schulz constructs a sequence of approximations $\bm{X}_{t}\approx\operatorname*{polar}(\bm{M})$ as follows:
@@
At each iteration, this rule effectively applies the cubic polynomial $p(x)=\frac{3}{2}x-\frac{1}{2}x^{3}$ to each singular value of $\bm{X}_{t}$. The scalar fixed-point iteration $x_{t+1}=p(x_{t})$ converges to $\operatorname*{sign}(x_{0})$ as $t\to\infty$, provided $|x_{0}|\leq 1$. As a result, the matrix iteration satisfies $\lim\limits_{t\to\infty}\bm{X}_{t}=\bm{U}\bm{V}^{\top}=\operatorname*{polar}(\bm{X}_{0})$. Higher-degree versions of Newton-Schulz follow the same principle. For example, the degree-5 polynomial $p(x)=(15x-10x^{3}+3x^{5})/8$ converges even faster. The Newton-Schulz iterations converge super-exponentially when $\bm{X}_{t}$ is sufficiently close to $\operatorname*{polar}(\bm{M})$, but they suffer from slow initial convergence; when $\bm{X}_{0}$ is far from $\operatorname*{polar}(\bm{M})$, the approximation improves slowly over the first few iterations.
The Jordan and You methods.
In Muon, high accuracy approximations to $\operatorname*{polar}(\bm{M})$ are usually not necessary. The primary goal is instead to compute a coarse approximation in as few iterations as possible. To accelerate convergence in the low-accuracy regime, Jordan recently proposed a fixed-point iteration based on the polynomial $p(x)=3.4445x-4.7750x^{3}+2.0315x^{5}$ [22], which was found using a heuristic numerical search. Unlike Newton-Schulz, the scheme that Jordan proposed does not converge to $\operatorname*{polar}(\bm{M})$. Instead, it plateaus at an error of $\approx 0.3$. However, it reaches this level of accuracy rapidly. As a result, when the number of iterations is smaller than about $10$, Jordan’s method outperforms the Newton-Schulz iteration. Building on this idea, You [9] proposed a method that applies six different polynomial updates in succession, which were again found by heuristic search. This method achieves better accuracy than Jordan’s but still fails to converge.
@@ 1.3   Contributions
We present Polar Express, an iterative method for approximating $\operatorname*{polar}(\bm{M})$. Our method dynamically adapts the polynomial update rule at each iteration, prioritizing rapid progress in the initial stage and high accuracy in the later stage. Polar Express constructs polynomials $p_{1},\ldots,p_{T}$ so that the resulting composition is the optimal approximation to the sign function with respect to the supremum ($L^{\infty}$) norm (Theorem 4.1). By iteratively applying these polynomials to $\bm{M}$, Polar Express computes an approximation to $\operatorname*{polar}(\bm{M})$ that is optimal in the worst-case at every iteration. Our method converges to $\operatorname*{polar}(\bm{M})$ super-exponentially (Theorem 4.3), and it quickly reaches a good approximation within just five to ten iterations. This early-stage acceleration is especially valuable in deep learning applications, where runtime efficiency takes precedence over high accuracy. In contrast, classical methods like Newton-Schulz suffer from a slow initial convergence, while recent heuristic proposals [22, 9] fail to converge. Our method is efficient to run on GPUs, using only a few matrix-matrix products per iteration. [^1] In Appendices E and F, we describe two further algorithmic ideas that can be incorporated into Polar Express. They are not used in our Muon experiments but they may be beneficial in other settings, and we believe they merit further study.
We give an explicit instantiation of Polar Express in Algorithm 1, which incorporates minor modifications to make it compatible with half-precision arithmetic (see Section 4.4). Algorithm 1 can be used as a drop-in replacement for previous methods. In numerical experiments, Polar Express outperforms previous methods on synthetic matrices and gradient matrices from a GPT-2 transformer (Figure 4). We demonstrate the effectiveness of using Polar Express within the Muon optimizer in Figure 1, showing that it consistently improves the training of GPT-2 language models on 1 billion tokens of the FineWeb dataset [2].
@@
For a polynomial $p\in\mathbb{P}_{d}^{\operatorname*{odd}}$ and a matrix $\bm{M}$ with rank reduced SVD given by $\bm{M}=\bm{U}\bm{\Sigma}\bm{V}^{\mathsf{T}}$ and positive singular values $\sigma_{1}\geq\cdots\geq\sigma_{\operatorname*{rank}(\bm{M})}>0$, we define $p(\bm{M}):=\bm{U}p(\bm{\Sigma})\bm{V}^{\mathsf{T}}$, where $p(\bm{\Sigma})$ is the diagonal matrix with diagonal entries $p(\sigma_{i})$ for $i=1,\ldots,\operatorname*{rank}(\bm{M})$.
@@ 2   Related Work
Newton-Schulz and polynomial Padé methods.
The earliest methods in the literature are polynomial iterations like (2). Several nearly simultaneous papers introduced the family of polynomial Padé iterations, comprising Newton-Schulz and its higher-degree analogues [27, 6, 18, 29]. These higher-degree methods are also sometimes called “Newton-Schulz”; when doing so, we will specify the degree for clarity. In these methods, each iteration refines the current approximation $\bm{X}_{t}$ by applying a low-degree odd matrix polynomial, where any odd monomial $x\mapsto x^{2q+1}$ is defined for rectangular matrices by the formula $\bm{X}_{t}\mapsto\bm{X}_{t}\left(\bm{X}_{t}^{\top}\bm{X}_{t}\right)^{q}$. Our Polar Express method also takes this form, though unlike Newton-Schulz, it changes the polynomial at each iteration.
The polynomials used in Padé methods are chosen to match the value and first few derivatives of $\operatorname*{sign}(x)$ at the points $x=\pm 1$. For instance, the update rule of the third method in this family is defined by $p(x)=\frac{1}{16}\left(35x-35x^{3}+21x^{5}-5x^{7}\right)$, which is the unique degree-7 polynomial satisfying $p(\pm 1)=\pm 1$ and $p^{\prime}(\pm 1)=p^{\prime\prime}(\pm 1)=p^{\prime\prime\prime}(\pm 1)=0$. These methods converge so long as all singular values of $\bm{X}_{0}$ lie in $(0,1]$, a condition guaranteed by the initialization of (2). Furthermore, the order of convergence of the degree $12q+$ 1 method is $q+1$ [6]. In particular, the Newton-Schulz method ($q=1$) converges quadratically.
Newton’s method and rational Padé.
In the numerical analysis literature, polynomial methods were succeeded by rational iterations like Newton’s method [18], defined as follows [^2] Our description of Newton’s method and other rational methods assumes square non-singular $\bm{M}$. Non-square problems can be reduced to the square case by an initial QR decomposition, but this is not an option for purely polynomial methods like ours.:
@@
Newton’s method also converges quadratically. Like Newton-Schulz, it works because the rational function $r(x)=\frac{1}{2}(x+x^{-1})$ has a stable fixed point at $1$; unlike for Newton-Schulz, this point is a global attractor for the whole positive real line. At first glance, Newton’s method has nothing to do with the Padé iterations discussed above. However, after a change of variables $\bm{Y}_{t}=\bm{X}_{t}^{-1}$, it can be reinterpreted as $\bm{Y}_{t+1}=2\bm{Y}_{t}(\bm{I}+\bm{Y}_{t}^{\top}\bm{Y}_{t})^{-1}$, which is sometimes called inverse Newton. Observing that $r(x)=\frac{2x}{1+x^{2}}$ satisfies $r(\pm 1)=\pm 1$ and $r^{\prime}(\pm 1)=0$, we see that (inverse) Newton is also a Padé method, though a rational rather than polynomial one. In fact, given a odd degree $12q_{n}+$ 1 for the numerator and an even degree $2q_{d}$ for the denominator, there is a unique rational function that matches the value and first $q_{n}+q_{d}$ derivatives of $\operatorname*{sign}(x)$ at $x=\pm 1$. This directly yields a Padé method for computing $\operatorname*{polar}(\bm{M})$ whose order of convergence is $q_{n}+q_{d}+1$. For instance, $r(x)=\frac{3x+x^{3}}{1+3x^{2}}$ is called Halley’s method, which converges cubically. When $q_{d}=0$, we recover the polynomial Padé methods.
@@
There are two main weakness of Newton’s method and the Padé iterations: slow convergence in the initial phase and the need to compute explicit inverses. To accelerate initial convergence, Higham popularized the technique of rescaling the matrix after every Newton iteration [18]. Intuitively, rescaling $\bm{X}_{t}$ so that $\sigma_{\max}=1/\sigma_{\min}$ centers the spectrum around $1$, where convergence is fastest. Several easily-computable choices of scaling factor exist to accomplish this approximately. Note that this rescaling scheme would fail for Newton-Schulz, which likewise suffers from slow initial convergence but which would diverge if $\sigma_{\max}\gg 1$.
Computing matrix inverses is difficult to parallelize and to implement stably in low precision arithmetic. However, a trick was developed for stably computing many rational methods without explicit inverses; QR decompositions can be used instead [32, 49]. Applying this trick to Halley’s method and combining with a special rescaling scheme yields the QDWH (QR-based dynamically weighted Halley) method, which converges in just six iterations for any reasonably conditioned matrix [32].
Adaptive rational methods from optimal approximations.